OpenStudy (anonymous):

10.03 Simplifying Radical Expressions 7 questions? Please help! :)

OpenStudy (anonymous):

1)Which expression is the simplified form of the square root of 150b^3? 2)Simplify the square root of 50x^6. 3)Which number = the square root of 100 (the answer is 50, right?). 4)Simplify the square root of 48. 5)What expression is the simplified form of the square root of y^4? 6)Which of the following equals 52?(it's 10, correct?) 7)Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, 7 to the square root of a can be written as 7sqrt(a).

OpenStudy (anonymous):

Any help is greatly appreciated! :)

OpenStudy (anonymous):

answer 1 \[5b\sqrt{6b}\] answer 2 \[5x ^{3}\sqrt{2}\] answer 3 \[\sqrt{100}=10\] answer 4 \[4\sqrt{3}\] answer 5 \[y ^{2}\] answer 6 could't understand

OpenStudy (anonymous):

Oh, 6 is 5^2, my bad, and thanks so much!! :-D

OpenStudy (anonymous):

answer 6 5^2=25

OpenStudy (anonymous):

Thanks, much!

OpenStudy (anonymous):

Glad to help!! Happy thanksgiving!!!